# Inflation with Gauss-Bonnet Correction and Higgs Potential

### v1

@@ Abstract
We investigate the cosmological inflation for the Einstein-Hilbert action plus the Higgs potential function and the Gauss-Bonnet term coupled with the Higgs scalar field through a dilaton-like coupling. Then, using the Friedmann-Lemaître-Robertson-Walker metric and considering the appropriate slow-roll parameters, we derive the necessary equations of motion. In the proposed model, since the $e$-folding integral cannot be easily solved analytically, we first utilize a well-known Taylor expansion. Then, with a certain range of values derived for the model parameters, utilizing a mixture of several diagrams and numerical analysis methods, we obtain results for the tensor-to-scalar ratio and the scalar spectral index that are in good agreement with observational data within the acceptable range of the $e$-folding values. Also, in the absence of the Gauss-Bonnet term, we find that the inflationary observables are roughly the same as the predictions of the chaotic inflation model.
@@ 1 Introduction
The cosmic inflation theory has been agreed upon by many cosmologists as a suitable way to explain fundamental problems of the standard model of cosmology, such as the flatness problem and the horizon problem in the early universe. For this purpose, the inflation generated based on a canonical scalar field with minimal coupling is one of the simplest methods available, wherein the scalar field slowly rolls during the inflation. In this type of model, the scalar field is called inflaton because it is the main cause of inflation [1, 2, 3, 4, 5, 6, 7]. According to this model, the quantum fluctuations of the inflaton scalar field affect the curvature perturbations, and with the growth of these perturbations, cosmic structures such as galaxies and stars are formed over time. To check the correctness of such models, the measurement of temperature anisotropy in the cosmic microwave background (CMB) can be used. The latest CMB data obtained from the Planck satellite show serious limitations on the values of the scalar spectral index of curvature perturbations ($n_{\rm S}$) and the tensor-to-scalar ratio ($r$). Such limitations indicate that there is not much freedom in choosing the parameters of the subject, including the type of inflaton and its potential [8, 9, 10].
@@
At present, the only scalar field available in the set of fundamental particles is the field corresponding to the Higgs boson, which appears to be a suitable candidate for the inflaton, see, e.g., Refs. [7, 12, 13, 14, 15, 16, 17, 18, 19, 20]. In particular, with the discovery of Higgs features in the large hadron collider, which shows that the Higgs boson model has significant uncertainties in its parameters compared to some of the models used to develop the standard model. However, if the Higgs scalar field enters the inflation directly, it will be incompatible with the standard model of particle physics. To solve this problem, it has been proposed to add a non-minimal coupling to the gravity term, which has been studied the most, see, e.g., Refs. [12, 21, 19, 22]. By considering the non-minimal coupling between gravity and the Higgs scalar field, it is possible to remedy the aforementioned inconsistency problem and stabilize inflation, but new problems are encountered at the same time. Indeed, the price given for the desirability of this model is the unusual magnitude of the non-minimal coupling, approximately $10^{4}$. Such a (meta-)stability of the vacuum in the Higgs inflation model is still under discussion [23, 24]. On the other hand, due to the presence of the coupling term with a large coupling constant, the unitarity bound of the theory is violated because the energy scale in this inflation model is larger than or close to the cutoff value [22].
@@
The Gauss-Bonnet term is the second term of the Lovelock Lagrangian. It is also induced from the superstring theory and a subclass of the Horndeski theory. However, including the Gauss-Bonnet term alone in a four-dimensional Lagrangian of gravitation does not affect the equations of motion, see, e.g., Refs. [26, 27, 28]. Actually, it is topologically invariant in $4$-dimensions, and hence non-dynamical in $4$-dimensions. Nonetheless, its effect can appear when coupled non-minimally with a scalar field function, which can then play as an effective potential in causing inflation. Of course, it creates violent negative instabilities for tensor perturbations around a de Sitter background on small scales [9, 10, 11, 29, 30, 31].
The inflation resulting from the coupling with the Gauss-Bonnet term has been studied in Refs. [8, 9, 10, 11, 24, 25, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54]. For simplicity in calculations and analysis, in a number of references such as [8, 9, 54], the potential function and the coupling function ($V(\phi)$ and $\xi(\phi)$ as functions of the inflaton scalar field) are assumed so that the relaxation condition $V(\phi)\xi(\phi)$ is a constant. In these references, the focus is on two specific models, namely ($V(\phi)=V_{0}e^{-\lambda\phi}$ and $\xi(\phi)=\xi_{0}e^{\lambda\phi}$ with $V_{0}$ and $\xi_{0}$ as constants) and ($V(\phi)=V_{0}\phi^{n}$ and $\xi(\phi)=\xi_{0}\phi^{-n}$), wherein the results of calculations lead to the reduction of the tensor-to-scalar ratio.
In this study, we focus on breaking such a relaxation condition to check whether the obtained results are consistent with the Planck observational data. Moreover, we consider the inflaton scalar field to be a Higgs scalar field with a Higgs potential. However within inflation, since the field is much larger than the expected vacuum value, the potential can be approximated to a quartic form, $V_{0}\phi^{4}$. On the other hand, in this work, we consider the coupling function as dilaton-like, i.e., $\xi(\phi)=\xi_{0}e^{-\lambda\phi}$, and investigate the scalar spectral index and the tensor-to-scalar ratio, and compare the results with the most recent constraints by the Planck collaboration [55], i.e.,
@@
$r<0.10\quad\left({\rm at}\ 95\%\,{\rm CL,PlanckTT+lowE+lensing}\right)$ (1)
and the joint Planck, BK1515 and BAO data in further constraint tightened the upper limit on r as
$r<0.056\quad{\rm at}\ 95\%\,{\rm CL}.$ (2)
@@ 3 The Proposed Model
Now, in the presence of the Gauss-Bonnet term, to proceed and compare the obtained results with the Planck observational data, the functionality of $\phi$ should be specified and then inserted into relations (42) and (43). In this regard, a routine procedure is to first acquire the inflaton scalar field as a function of the $e$-folding number. Thereupon, by inserting the best value of $e$-folding number (around 607060-70), the obtained results can be checked easily. Hence, we must solve relation (22) by knowing the inflaton scalar field at the end of inflation and then obtain the $e$-folding number as a function of the inflaton scalar field. Next, by deriving the inverse function, we can achieve the inflaton scalar field as a function of the $e$-folding number.
@@
Now, by substituting the fitted Eqs. (60)-(63) into relations (42) and (43), we obtain the relations of $r$ and $n_{\rm S}$ with respect to the $e$-folding number and the model parameters. Then, to get the best results comparable with the data (1) and (2), using the obtained relations, we have drawn $r$ versus $n_{\rm S}$ in Fig. 6 for a few acceptable $e$-folding numbers.
@@
This figure shows that for two sets of the model parameters, i.e., $\alpha=0.01$ with $\lambda=0.48$ and $\alpha=0.02$ with $\lambda=0.57$, the obtained results for the inflationary observables are in good agreement with the observational data. Our investigations show that to have acceptable values for the inflationary observables consistent with the Planck observational data, the best model parameters in the range $N\in(50,70]$ should be in the range $\alpha\in(0,0.02]$ and slightly more with (almost) the corresponding minimum value of $\lambda$ specified by constraint (53) for each $\alpha$. The observable quantities are sensitive to the parameter $\alpha$, and when $\alpha$ increases from $0.02$, the results begin to deviate from the Planck observational data, and thus there is a small range for values of the model parameters to provide acceptable results.
@@ 4 Conclusions
Finally, for the proposed model, with a certain range of values derived for the model parameters, we have obtained results for the tensor-to-scalar ratio and the scalar spectral index that are in good agreement with observational data within the acceptable range of the $e$-folding values. However, there is not much choice for the values of the model parameters. Of course, one of the key points of this work is to achieve results through analytical analysis methods. Also, in the absence of the Gauss-Bonnet term, we find that the inflationary observables are roughly the same as the predictions of the chaotic inflation model. However, in this case, the value obtained for the tensor-to-scalar ratio is not consistent with the Planck observable data, while the very acceptable values obtained for it in the presence of the Gauss-Bonnet term indicate the significant effect of including the Gauss-Bonnet correction. In addition, the results of the model show that the speed of gravitational waves during inflation is greater than one (i.e., the speed of light) but very close to it, and its very slight difference from one is due to the presence of the Gauss-Bonnet correction.

### v2

@@ Abstract
We investigate the cosmological inflation for the Einstein-Hilbert action plus the Higgs potential function and the Gauss-Bonnet term coupled with the Higgs scalar field through a dilaton-like coupling. Then, using the Friedmann-Lemaître-Robertson-Walker metric and considering the appropriate slow-roll parameters, we derive the necessary equations of motion. In the proposed model, since the $e$-folding integral cannot be easily solved analytically, we first utilize a well-known Taylor expansion. Then, with a certain range of values derived for the model parameters, utilizing a mixture of several diagrams and numerical analysis methods, we obtain results for the tensor-to-scalar ratio and the scalar spectral index that are in good agreement with the latest observational data, particularly from ACT DR6, within the acceptable range of the $e$-folding values. Also, in the absence of the Gauss-Bonnet term, we find that the inflationary observables are roughly the same as the predictions of the chaotic inflation model.
@@ 1 Introduction
The cosmic inflation theory has been agreed upon by many cosmologists as a suitable way to explain fundamental problems of the standard model of cosmology, such as the flatness problem and the horizon problem in the early universe. For this purpose, the inflation generated based on a canonical scalar field with minimal coupling is one of the simplest methods available, wherein the scalar field slowly rolls during the inflation. In this type of model, the scalar field is called inflaton because it is the main cause of inflation [1][8]. According to this model, the quantum fluctuations of the inflaton scalar field affect the curvature perturbations, and with the growth of these perturbations, cosmic structures such as galaxies and stars are formed over time. To check the correctness of such models, the measurement of temperature anisotropy in the cosmic microwave background (CMB) can be used. The latest CMB data obtained from the Planck satellite and the recent sixth data release (DR6) of the Atacama Cosmology Telescope (ACT) collaboration show serious limitations on the values of the scalar spectral index of curvature perturbations, $n_{\rm S}$, and the tensor-to-scalar ratio, $r$. Such limitations indicate that there is not much freedom in choosing the parameters of the subject, including the type of inflaton and its potential [9][11].
@@
At present, the only scalar field available in the set of fundamental particles is the field corresponding to the Higgs boson, which appears to be a suitable candidate for the inflaton, see, e.g., Refs. [8, 13][21]. In particular, with the discovery of Higgs features in the large hadron collider, which shows that the Higgs boson model has significant uncertainties in its parameters compared to some of the models used to develop the standard model. However, if the Higgs scalar field enters the inflation directly, it will be incompatible with the standard model of particle physics. To solve this problem, it has been proposed to add a non-minimal coupling to the gravity term, which has been studied the most, see, e.g., Refs. [13, 20, 22, 23]. By considering the non-minimal coupling between gravity and the Higgs scalar field, it is possible to remedy the aforementioned inconsistency problem and stabilize inflation, but new problems are encountered at the same time. Indeed, the price given for the desirability of this model is the unusual magnitude of the non-minimal coupling, approximately $10^{4}$. Such a (meta-)stability of the vacuum in the Higgs inflation model is still under discussion [24, 25]. On the other hand, due to the presence of the coupling term with a large coupling constant, the unitarity bound of the theory is violated because the energy scale in this inflation model is larger than or close to the cutoff value [23]. Recently, a Bayesian analysis of the Starobinsky and Higgs inflationary models has been performed in Ref. [26].
@@
The Gauss-Bonnet term is the second term of the Lovelock Lagrangian. It is also induced from the superstring theory and a subclass of the Horndeski theory [28][30]. However, including the Gauss-Bonnet term alone in a four-dimensional Lagrangian of gravitation does not affect the equations of motion, see, e.g., Refs. [31][33]. Actually, it is topologically invariant in $4$-dimensions, and hence non-dynamical in $4$-dimensions. Nonetheless, its effect can appear when coupled non-minimally with a scalar field function, which can then play as an effective potential in causing inflation. Of course, it creates violent negative instabilities for tensor perturbations around a de Sitter background on small scales [10][12, 35][37].
The inflation resulting from the coupling with the Gauss-Bonnet term has been studied in Refs. [9][12, 25, 27, 34][62]. For simplicity in calculations and analysis, in a number of references such as [9, 10, 60], the potential function and the coupling function ($V(\phi)$ and $\xi(\phi)$ as functions of the inflaton scalar field) are assumed so that the relaxation condition $V(\phi)\xi(\phi)$ is a constant. In these references, the focus is on two specific models, namely ($V(\phi)=V_{0}e^{-\lambda\phi}$ and $\xi(\phi)=\xi_{0}e^{\lambda\phi}$ with $V_{0}$ and $\xi_{0}$ as constants) and ($V(\phi)=V_{0}\phi^{n}$ and $\xi(\phi)=\xi_{0}\phi^{-n}$), wherein the results of calculations lead to the reduction of the tensor-to-scalar ratio.
In this study, we focus on breaking such a relaxation condition to check whether the obtained results are consistent with the observational data. Moreover, we consider the inflaton scalar field to be a Higgs scalar field with a Higgs potential. However within inflation, since the field is much larger than the expected vacuum value, the potential can be approximated to a quartic form, $V_{0}\phi^{4}$. On the other hand, in this work, we consider the coupling function as dilaton-like, i.e., $\xi(\phi)=\xi_{0}e^{-\lambda\phi}$, and investigate the scalar spectral index and the tensor-to-scalar ratio, and compare the results with the recent constraints by the Planck collaboration [63, 64]
@@
$r<0.10\quad\left({\rm at}\ 95\%\,{\rm CL,PlanckTT+lowE+lensing}\right)\!,$ (1)
the joint Planck, BK1515 and BAO data, which places a further constraint on the upper limit of $r$,
$r<0.056\quad{\rm at}\ 95\%\,{\rm CL},$ (2)
and the most recent P-ACT-LB DR6 [65, 66] data, which reports a higher $n_{\rm S}$ value and still tightens the upper limit of $r$,
$n_{\rm S}=0.9743\pm 0.0034,$ (3)
$r<0.038\quad{\rm at}\ 95\%\,{\rm CL}.$ (4)
@@ 3 The Proposed Model
Now, in the presence of the Gauss-Bonnet term, to proceed and compare the obtained results with the observational data, the functionality of $\phi$ should be specified and then inserted into relations (44) and (45). In this regard, a routine procedure is to first acquire the inflaton scalar field as a function of the $e$-folding number. Thereupon, by inserting the best value of $e$-folding number (around 607060-70), the obtained results can be checked easily. Hence, we must solve relation (24) by knowing the inflaton scalar field at the end of inflation and then obtain the $e$-folding number as a function of the inflaton scalar field. Next, by deriving the inverse function, we can achieve the inflaton scalar field as a function of the $e$-folding number.
@@
Now, by substituting the fitted Eqs. (62)-(65) into relations (44) and (45), we obtain the relations of $r$ and $n_{\rm S}$ with respect to the $e$-folding number and the model parameters. Then, to get the best results comparable with the data (1), (2) and (3), using the obtained relations, we have drawn $r$ versus $n_{\rm S}$ in Fig. 6 for a few acceptable $e$-folding numbers.
@@
This figure shows that for two sets of the model parameters, i.e., $\alpha=0.01$ with $\lambda=0.48$ and $\alpha=0.02$ with $\lambda=0.57$, the obtained results for the inflationary observables are in good agreement with the observational data. Our investigations show that to have acceptable values for the inflationary observables consistent with the observational data, the best model parameters in the range $N\in(50,70]$ should be in the range $\alpha\in(0,0.02]$ and slightly more with (almost) the corresponding minimum value of $\lambda$ specified by constraint (55) for each $\alpha$. The observable quantities are sensitive to the parameter $\alpha$, and when $\alpha$ increases from $0.02$, the results begin to deviate from the observational data, and thus there is a small range for values of the model parameters to provide acceptable results.
@@ 4 Conclusions
Finally, for the proposed model, with a certain range of values derived for the model parameters, we have obtained results for the tensor-to-scalar ratio and the scalar spectral index that are in good agreement with the latest observational data, particularly from ACT DR6, within the acceptable range of the $e$-folding values. However, there is not much choice for the values of the model parameters. Of course, one of the key points of this work is to achieve results through analytical analysis methods. Also, in the absence of the Gauss-Bonnet term, we find that the inflationary observables are roughly the same as the predictions of the chaotic inflation model. However, in this case, the value obtained for the tensor-to-scalar ratio is not consistent with the Planck observable data, while the very acceptable values obtained for it in the presence of the Gauss-Bonnet term indicate the significant effect of including the Gauss-Bonnet correction. In addition, the results of the model show that the speed of gravitational waves during inflation is greater than one (i.e., the speed of light) but very close to it, and its very slight difference from one is due to the presence of the Gauss-Bonnet correction.